# Measurement-induced phase transition in periodically driven free-fermionic systems

### v1

@@ IV.2 Numerical results
Given that the steady entanglement entropy growth is logarithmic in a small $\gamma$ regime, we also use the similar BKT scaling ansatz to obtain the critical $\gamma_{c}$ values for different $\epsilon$ and different values of $T$. Figure. 7 (lower panel) shows the data collapse of steady state entanglement entropy for different $L$ and $\gamma$, which confirms that the transition belongs to the BKT universality class and allows us to evaluate the critical $\gamma_{c}$. We use the cost function minimization technique to obtain the $\gamma_{c}$ (see Appendix. B). Inset shows the critical $\gamma_{c}$ values for different $\epsilon$. Our results suggest the critical $\gamma_{c}$ decreases with decreasing $\epsilon$ and approaches zero in the $\epsilon\to 0$ limit, as also predicted by our previous RG study for high-frequency limit. The best fit suggests $\gamma_{c}\sim(\epsilon/2J)^{\nu}$, were $\nu\simeq 0.91$, slightly smaller than $1$. Note that $\nu=1$ was predicted previously in sec. IV.1 for the high-frequency limit. Figure. 8 shows the contour plot of the effective central charge $C(\gamma)$ for different values of $\gamma$ and $\epsilon$, and dots correspond to the $\gamma_{c}$ and the dashed line corresponds to the best-fit line. Also, important to point out, given our $T=0.5$ data are almost indistinguishable from the no-drive data with $J_{eff}=\epsilon/2$, one will obtain $\nu=1$ for $T=0.5$. $\nu<1$ for $T=5$ also signifies that with increasing $T$, $\gamma_{c}$ will increase for a given $\epsilon$; it will promote the critical phase more. Moreover, the data for $T=5$ also suggests that for a symmetric drive with zero means, i.e., for $\epsilon=0$, $\gamma_{c}\to 0$, implies no transition, always area-law phase. Though our analytical results suggest the same for the high-frequency regime; it is not at all obvious that this phenomenon persists even for large $T$. Especially given that there are examples where disorder-induced localization-delocalization transition has been observed in the large $T$ limit for symmetric drive mean around zero Sarkar et al. (2021, 2022). Hence, in the next section, we investigate the $\epsilon=0$ limit even more carefully.
@@
Figure 10: Upper panel: Scaling of the steady state entanglement entropy with system size $L$ for different $T$, but fixed $\gamma=0.05$ and $\epsilon=0$. Initially, it shows logarithmic growth with $L$ and then tends to saturate. Dashed-line corresponds to no-drive $J=1/2$ results for $\gamma=0.05$. Lower panel: plot of $L_{c}$ vs. $T$, $L_{c}$ is estimated using Eq. (16), circles are for square pulse protocol, and triangles are for sinusoidal protocol. Inset shows the best fit of the entanglement entropy assuming the functional form given by Eq. (16) for square pulse protocol.
@@
IV.3 Symmetric drive with zero mean: $\epsilon=0$ limit
@@
Square pulse drive: Here, we focus on the $\epsilon=0$ limit. First, the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial Neel state for different values of time period $T$ and measurement strengths $\gamma,$ is studied. We find that $\overline{S}$ increases with time initially and then saturates to a steady state value. From Fig. 9 upper panel, we see, for a fixed driving period $T=20$, if we increase the measurement strength $\gamma$, the steady state entanglement saturation value is decreased; similar to our previous results, the entanglement spread gets suppressed because of increasing measurement strength. On the other hand, it is shown in the lower panel of Fig. 9 that, for a fixed measurement strength $\gamma=0.05$, the entanglement saturation value increases with the increase of the driving time-period $T$. This suggests, in general, that the low driving frequency promotes the depinning of the particles, and the entanglement entropy saturation values approach the no-drive result with increasing $T$. We also noticed there is a critical driving period $T_{c}$, such that for $T>T_{c}$, the results are almost indistinguishable from those of the no-drive scenario. However, this $T_{c}$ typically increases with $L$. The fact that $T>T_{c}$ resembles a no-drive situation has also been observed for our $2\times 2$ Toy model previously.
@@
Next, we immediately ask the following question: How does the steady-state entanglement entropy change with the subsystem size for different $T$? The existence of a $T_{c}$ for each $L$ might be a tempting scenario to conjecture that in the very large $T$ limit, $\epsilon=0$ results should resemble no-drive results and should display an entanglement transition. However, one needs to take the large $T$ and large $L$ limits carefully. The order of these two limits doesn’t commute. If one takes a large $T$ limit first, then any finite $L$ (no matter how large it is), $T_{c}$ has been achieved, and hence, one will observe no-drive results, i.e., the presence of the measurement-induced transition. On the other hand, the correct limit is to take the large $L$ limit first, then increase the $T$. In that case, the outcome will be different, and we further try to argue that in this limit, there will be no transition.
@@
Given for each $L$, a critical $T_{c}$ exists, and if one chooses a reasonably large $T$, it will turn out to be a $T_{c}$ for some system size $L=L_{1}$. Hence, for that particular $T$, till $L<L_{1}$, the saturation value will resemble the no-drive results, which implies it will show logarithmic scaling for $L<L_{1}$ for small $\gamma$, but for $L>L_{1}$, the steady entanglement data vs $L$ will certainly show a bending because if $T<T_{c}$, the entanglement saturation value is smaller than the $T\geq T_{c}$ value. We precisely see that in Fig. 10 upper panel. We find for $\gamma=0.05$ and $T=5$, and $8$, the steady state entanglement entropy is almost constant for all $L$, which suggests an area law scaling with the system size. On the other hand, as we increase $T$, we see an initial logarithmic scaling of entanglement entropy with system size till $L<L_{1}$, followed by bending trend, finally after a critical $L=L_{c}$, entanglement starts showing area law behavior. The dashed line corresponds to the no-drive results for $J=1/2$ and $\gamma=0.05$. Roughly, till $L\simeq 20$, this no-drive results resembles $T=20$ data, beyond $L=20$, the no-drive results keep increasing with $L$, while the $T=20$ data starts bending (see appendix. A for $T=100$ results). To estimate the $L_{c}$, we fit the steady entanglement entropy vs $L$ data with a two parameters ($S_{sat}$ and $L_{c}$) fitting function $f(L)$, i.e.,
@@
Sinusoidal drive: To become more certain about our findings for the $\epsilon=0$ limit, we also study the effect of sinusoidal driving, where $J(t)=J\sin(2\pi t/T)$. We find that even for the sinusoidal drive protocol, the area-law entanglement phase sustains in the large $L$ limit. Growth of trajectory averaged entanglement entropy for $T=20$ and $\gamma=0.05$ are shown in Fig. 11 for different $L$. We find that $L=400$ and $550$ data are almost indistinguishable, though for small $L$, the saturation values are different. The inset of Fig. 11 shows the variation of the steady state entanglement entropy with $L$ for different $T$, and it clearly shows the evidence of the area law in the large $L$ limit, as already observed for the square pulse case. We also estimate the $L_{c}$ using the fitting function $f(L)$ (see Eq. (16)) and compare it with square pulse results in Fig. 10 lower panel. We find that $L_{c}$ is even smaller for sinusoidal drive compared to the square pulse.
Figure 11: Stroboscopical growth of trajectory averaged entanglement entropy for different $L$ for the sinusoidal driving protocol for $T=20$ and $\gamma=0.05$(main panel). Inset shows the system size scaling of the steady state trajectory averaged entanglement entropy for the different driving periods $T$. Data presented here are for $\gamma=0.05$.
@@ V Conclusions
In this work, we investigate the effect of periodic drive in a continuously monitored free-fermionic lattice model. It is well-known that in the absence of the drive, such systems undergo a measurement-induced entanglement phase transition from critical (logarithmic) to area law. Our main finding is that periodically driven non-interacting systems also go through a measurement-induced entanglement phase transition, and the universality class of this transition also seems to be preserved, which belongs to the BKT universality class. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that the critical measurement strength (transition point) increases with the time period. Moreover, we have also investigated a drive protocol in which the hopping amplitudes are varied completely symmetrically around zero in the form of square-pulse and sinusoidal drive and found that such a drive can’t cause a phase transition independent of any frequency regime; it always promotes an area-law entanglement phase in the thermodynamic limit. This result is a bit surprising, especially given that disorder-induced similar-driven systems tend to favor de-localization in the low-frequency regime Sarkar et al. (2021). Our results suggest several interesting directions for future work. The role of disorders, interactions in the monitored Floquet system, and their relationship to Floquet-MBL Sierant et al. (2023) remain a challenging problem.

### v2

@@ IV.2 Numerical results
Given that the steady entanglement entropy growth is logarithmic in a small $\gamma$ regime, we also use a similar BKT scaling ansatz to obtain the critical $\gamma_{c}$ values for different $\epsilon$ and $T=5$. Figure. 7 (lower panel) shows the data collapse of steady state entanglement entropy for different $L$ and $\gamma$, which allows us to evaluate the critical $\gamma_{c}$. We use the cost function minimization technique to obtain the $\gamma_{c}$ (see Appendix. B). The inset shows the critical $\gamma_{c}$ values for different $\epsilon$. Our results suggest the critical $\gamma_{c}$ decreases with decreasing $\epsilon$ and approaches zero in the $\epsilon\to 0$ limit, as also predicted by our previous RG study for the high-frequency limit. Figure. 8 shows the contour plot of the effective central charge $c(\gamma)$ (which is obtained by fitting the steady-state entanglement data with $L$) for different values of $\gamma$ and $\epsilon$, and dots correspond to the $\gamma_{c}$. We discuss $\epsilon/2J=1$ (no-drive) results separately in the Appendix D. However, obtaining $\gamma_{c}$ accurately for large $T$ is extremely difficult. It turns out that for a given large $T$, there are system sizes $L<L_{1}$, for which the steady-state entanglement entropy is the same as the no-drive scenario with $J_{\text{eff}}=|J-\epsilon|$. The effect of the drive kicks in for $L>L_{1}$, where $L_{1}$ increases with $T$. This makes it extremely difficult to access the system size regime required to obtain the proper scaling. However, we do a comparative study of the results for different $T$ in Sec. IV.2 later. Moreover, the data for $T=5$ also suggests that for a symmetric drive with zero means, i.e., for $\epsilon=0$, $\gamma_{c}\to 0$, implies no transition, always an area-law phase. Though our analytical results suggest the same for the high-frequency regime, it is not at all obvious that this phenomenon persists even for large $T$. Especially given that there are examples where disorder-induced localization-delocalization transition has been observed in the large $T$ limit for symmetric drive mean around zero Sarkar et al. (2021, 2022). Hence, in Section IV.3, we investigate the $\epsilon=0$ limit even more carefully.
Methodology 2:
Figure 9: $T=5$ data of entanglement entropy scaling with subsystem size $l$(upper panel) and finite-size data collapse of entanglement entropy for different subsystem size(lower panel) for $L=800$.
Previously, to compute the steady state entanglement, we had restricted ourselves to the sub-system size $l=L/2$ and then increased $L$ to obtain finite size scaling. In this section, we present the entanglement entropy data for which we fix the total system size to $L\simeq 10^{3}$, and vary the sub-system size $l$. Once again, the entanglement entropy can be computed using Eq. (7), and we present the data of entanglement growth with subsystem size in Fig.  9 upper panel for $T=5$ and $L=800$ for $\epsilon=3J/4$. To obtain the critical $\gamma_{c}$, we use the same cost function minimization technique satisfying the scaling law in the vicinity of the critical phase as,
$\overline{S(l,\gamma)}-\overline{S(l,\gamma_{c})}=F[(\gamma-\gamma_{c})(\ln(\frac{L}{\pi}\sin{\frac{\pi l}{L}}))^{2}].$ (18)
Fig.  9 lower panel shows the data collapse of the entanglement entropy for different subsystem size $l$, with $\gamma_{c}=0.05\pm 0.014$ calculated using the cost function minimization presented in the Appendix.  B in Fig. 20.
Local effective central charge:
Figure 10: Effective central charge $c^{eff}_{l}$ vs. $l$ plots for different $\gamma$ values for $T=5,\epsilon=3J/4$. Black circular points represent the maximum of the central charge and red star points represent the position of $l^{*}_{c}$ for which $c^{eff}_{l}$ is the maximum estimated from the fitting. The inset shows the best fit, which gives $l^{*}_{c}\sim\gamma^{-1.8}$. The maximum system size is taken as $L=1004$.
Here we compute subsystem-dependent effective central charge and analyze our data as was done in a recent Ref.  Starchl et al. (2024b) for the no-drive scenario. We calculate the effective central charge defined as
$c^{eff}_{l}=3\frac{\overline{S}_{l^{\prime}}-\overline{S}_{l}}{\ln{\frac{\tilde{l^{\prime}}}{\tilde{l}}}},$ (19)
where, $\tilde{l}=\frac{L}{\pi}\sin(\pi l/L)$ is the cord length and $l$ is the size of the subsystem varies from $1$ to $L/2$. For logarithmic growth of the entanglement entropy, this measure should be constant; for volume-law scaling, it should scale as $c^{eff}_{l}\sim l$, while area-law behavior leads to $c^{eff}_{l}=0$. To calculate $c^{eff}_{l}$ numerically, we choose $N_{l}$ values of $l_{i}$ with $l_{1}=1$ and $l_{N_{l}}=L/2$ such that the successive ratio of $\frac{\tilde{l^{\prime}}}{\tilde{l}}$ is approximately constant. In Ref. Starchl et al. (2024b), it was shown that the critical phase is bounded from below by $l\sim\gamma^{-1}$ and from above by $l\sim\gamma^{-2}$. If the effective central charge increases with system size and attains a maximum, then it slowly starts decreasing, which indicates a crossover to area law. In Fig. 10 we present the effective central charge $c^{eff}_{l}$ data for the driving period $T=5$. We see a very clear evidence of bending of $c^{eff}_{l}$ after a critical $l^{*}_{c}$ for $\gamma>\gamma_{c}=0.05\pm 0.014$ (black circles represent the maximum). We fit those $l^{*}_{c}$ and obtain the scaling $l^{*}_{c}\sim\gamma^{-1.8}$. Following the fitting results, we estimate the $l^{*}_{c}$ values for $\gamma\leq\gamma_{c}$, which is represented by the red star points. Then we compare these $l^{*}_{c}$ with actual simulated data for system size $1004$, and we get only a very mild signature of bending for $\gamma=0.04$ (which is still in the regime of the critical $\gamma_{c}$ with our estimated error-bar), but for $\gamma=0.03,0.02$ (they are $<\gamma_{c}$) the bending is not at all clearly evident for the mentioned system size. We can interpret this in two different ways, either $l^{*}_{c}$ scales faster than power-law ${}\gamma^{-1.8}$ in the case of small $\gamma$ limit, so that the $l^{*}_{c}$ (beyond which $c^{eff}_{l}$ will show the bending) is very large for small $\gamma$ values, which is not accessible numerically, or there exist a critical phase for $\gamma<\gamma_{c}$, in that case data will never bend. Results in the absence of periodic drive are also presented in the Appendix. D for our protocol for comparison.
Dependancy of steady state entanglement entropy on driving frequency:
Figure 11: Comparison of entanglement entropy growth for different driving periods $T$ with system size $L=1000$ for $\gamma=0.05$. The inset shows that the slope of the entanglement growth is increasing with increasing $T$ for $\epsilon=3J/4$.
Figure 12: Effective central charge $c^{eff}_{l}$ vs. $l$ plots for different $T$ values for $\gamma=0.1,\epsilon=3J/4$. Black circular points represent the maximum of the central charge. The inset shows that $l_{c}^{*}$ is increasing with increasing $T$. The maximum system size is taken as $L=1004$.
Figure 13: $c_{l}^{eff}$ vs. $l-l_{c}^{*}$ plot with $l>l_{c}^{*}$ for different $T$ values. $l_{c}^{*}$ represents the value of $l$ for which $c_{l}^{eff}$ shows a peak. Black dashed lines shows the fitting having the functional form $c_{l}^{eff}\sim e^{-(l-l_{c}^{*})/\chi}$. The inset shows that $\chi$ is increasing with increasing $T$. The maximum system size is taken as $L=1004$.
Next, we investigate the effect of $T$ in the steady state entanglement scaling with subsystem size for a fixed $\epsilon=3J/4$. As mentioned earlier, as we increase $T$, it becomes even harder to compute $\gamma_{c}$; hence, we instead study the steady state entanglement scaling vs subsystem size for a given $\gamma$ and for different $T$. Figure 11 shows that in the large $l$ limit, the slope of the trajectory-averaged long-time entanglement entropy increases with increasing driving period $T$. The inset shows the slope of the entanglement entropy for different $T$, where we calculate the slope by fitting the entanglement entropy data in the large $l$ limit for $\gamma=0.05$ and $\epsilon=3J/4$, for system size $L=1000$. If the BKT transition survives in the thermodynamics limit for $T=5$ (which was one of the potential outcomes of our previous analysis), we can now infer from these results that increasing $T$ favors the critical phase more, and if the BKT transition survives, then the critical measurement strength $\gamma_{c}$ will increase with increasing $T$.
Next, to understand the effect of the driving frequency on measurement-induced transition in more detail, we investigate a scenario where we fix $\gamma=0.1$ for which $T=5$ data of effective local central charge clearly shows bending in the large $l$ limit (suggesting area law scaling), and then increase $T$. We observe two things (see Fig. 12): 1) the $l^{*}_{c}$ (the sub-system size for which the $c^{eff}_{l}=c_{eff}^{max}$ is maximum) increases with $T$ (this $c_{eff}^{max}$ is a measure of the maximum growth rate of the steady state entanglement with sub-system size, which we have plotted in Fig. 1 along with the schematic of our protocol), 2) the bending of the $c^{eff}_{l}$ data for $l>l^{*}_{c}$ becomes much slower with increasing $T$. To quantify the bending rate of the data, we plot $c^{eff}_{l}$ vs $(l-l_{c}^{*})$ for $l>l^{*}_{c}$ in Fig. 13, and find that it decreases exponentially with $l$. We fit the data with the function $c_{l}^{eff}\sim e^{-(l-l_{c}^{*})/\chi}$ for $l>l_{c}^{*}$, where a length scale $\chi$ is a measure of how much the data has bend, and $\chi\to\infty$ limit should correspond to the absence of the bending of $c_{l}^{eff}$ data, and that implies $c_{l}^{eff}$ is constant beyond $l=l^{*}_{c}$. This will be a signature of the critical phase. The inset of Fig. 13 shows, as expected, $\chi$ increases with increasing $T$. However, it is not clear from the data whether $\xi$ approaches infinity for any finite $T=T^{*}$ or not. At least it is clear that for $T\leq 100$, the critical $\gamma_{c}$ (assuming BKT transition exists) is certainly $<0.1$. We also investigate the dependence of $T$ on the steady state entanglement fluctuation in Appendix F.
IV.3 Dependency of steady state entanglement entropy on asymmetry parameter $\epsilon$:
Here we investigate the role of $\epsilon$, the asymmetry parameter in the driving amplitude in measurement-induced transition. In the Fig. 14, we see that the entanglement entropy slope increases with increasing $\epsilon$ for a fixed $T=50$. This also indirectly implies that if the BKT transition persists in the thermodynamic limit, then $\gamma_{c}$ will also increase with $\epsilon$. Similar results were obtained previously for $T=5$ as well. Next, we study the consequences of symmetric drive.
Figure 14: Comparison of entanglement entropy growth for different $\epsilon$ values with system size $L=800$ for $\gamma=0.05,T=50$. The inset shows that the slope of the entanglement growth is increasing with increasing $\epsilon$ for $T=50$.
@@ IV.3 Dependency of steady state entanglement entropy on asymmetry parameter $\epsilon$:
Square pulse drive with zero mean: $\epsilon=0$ limit:
@@
Here, we focus on the $\epsilon=0$ limit. First, the time evolution of the entanglement entropy, averaged over many different trajectories, is studied starting from the initial Neel state for various values of the time period $T$ and measurement strengths $\gamma$. We find that $\overline{S}$ increases initially with time and then saturates to a steady-state value. From Fig. 15 upper panel, we see, for a fixed driving period $T=20$, if we increase the measurement strength $\gamma$, the steady state entanglement saturation value is decreased; similar to our previous results, the entanglement spread gets suppressed because of increasing measurement strength. On the other hand, it is shown in the lower panel of Fig. 15 that, for a fixed measurement strength $\gamma=0.05$, the entanglement saturation value increases with the increase of the driving time-period $T$. This suggests, in general, that the low driving frequency promotes the depinning of the particles, and the entanglement entropy saturation values approach the no-drive result with increasing $T$. We also noticed there is a critical driving period $T_{c}$, such that for $T>T_{c}$, the results are almost indistinguishable from those of the no-drive scenario. However, this $T_{c}$ typically increases with $L$. The fact that $T>T_{c}$ resembles a no-drive situation has also been observed for our $2\times 2$ Toy model previously.
@@
Next, we immediately ask the following question: How does the steady-state entanglement entropy change with the subsystem size for different $T$? The existence of a $T_{c}$ for each $L$ might be a tempting scenario to conjecture that in the very large $T$ limit, $\epsilon=0$ results should resemble no-drive results and should display an entanglement transition for finite-size systems. However, one needs to take the large $T$ and large $L$ limits carefully. The order of these two limits doesn’t commute. If one takes a large $T$ limit first, then for any finite $L$ (no matter how large it is), $T_{c}$ has been achieved, and hence, one will observe no-drive results. On the other hand, the correct limit is to take the large $L$ limit first, then increase the $T$. In that case, for a given $T$ and a given $\gamma$ up to a system size $L=L_{1}$, the finite $T$ data will be same as the no-drive results. However, for $L>L_{1}$, the result will deviate from the no-drive result since their $T_{c}$ has not been reached. Clearly if $T_{c}$ increases with $L$, taking $L\to\infty$ first indicates that there cannot be any $T_{c}$ at all, so the results will never match the no-drive case.
Figure 16: Upper panel: Scaling of the steady state entanglement entropy with system size $L$ for different $T$, but fixed $\gamma=0.05$ and $\epsilon=0$. Initially, it shows logarithmic growth with $L$ and then tends to saturate. Dashed-line corresponds to no-drive $J=1/2$ results for $\gamma=0.05$. Lower panel: plot of $L_{c}$ vs. $T$, $L_{c}$ is estimated using Eq. (20), circles are for square pulse protocol, and triangles are for sinusoidal protocol. The inset shows the best fit of the entanglement entropy assuming the functional form given by Eq. (20) for the square pulse protocol.
@@
We precisely see that in Fig. 16 upper panel. We find for $\gamma=0.05$ and $T=5$, and $8$, the steady state entanglement entropy is almost constant for all $L$, which suggests an area law scaling with the system size. On the other hand, as we increase $T$, we see an initial logarithmic scaling of entanglement entropy with system size till $L<L_{1}$, followed by bending trend, finally after a critical $L=L_{c}$, entanglement starts showing area law behavior. The dashed line corresponds to the no-drive results for $J=1/2$ and $\gamma=0.05$. Roughly, till $L\simeq 20$, this no-drive results resembles $T=20$ data, beyond $L=20$, the no-drive results keep increasing with $L$, while the $T=20$ data starts bending (see appendix. A for $T=100$ results). To estimate the $L_{c}$, we fit the steady entanglement entropy vs $L$ data with a two parameters ($S_{sat}$ and $L_{c}$) fitting function $f(L)$, i.e.,
@@
Sinusoidal drive: To become more certain about our findings for the $\epsilon=0$ limit, we also study the effect of sinusoidal driving, where $J(t)=J\sin(2\pi t/T)$. We find that even for the sinusoidal drive protocol, the area-law entanglement phase sustains in the large $L$ limit. Growth of trajectory averaged entanglement entropy for $T=20$ and $\gamma=0.05$ are shown in Fig. 17 for different $L$. We find that $L=400$ and $550$ data are almost indistinguishable, though for small $L$, the saturation values are different. The inset of Fig. 17 shows the variation of the steady state entanglement entropy with $L$ for different $T$, and it clearly shows the evidence of the area law in the large $L$ limit, as already observed for the square pulse case. We also estimate the $L_{c}$ using the fitting function $f(L)$ (see Eq. (20)) and compare it with square pulse results in Fig. 16 lower panel. We find that $L_{c}$ is smaller for the sinusoidal drive compared to the square pulse.
Figure 17: Stroboscopical growth of trajectory averaged entanglement entropy for different $L$ for the sinusoidal driving protocol for $T=20$ and $\gamma=0.05$(main panel). The inset shows the system size scaling of the steady state trajectory-averaged entanglement entropy for the different driving periods $T$. The data presented here are for $\gamma=0.05$.
@@ V Conclusions
In this work, we investigate the effect of a periodic drive in a continuously monitored free-fermionic lattice model. We find that in the high-frequency limit $T<1$, the results are almost indistinguishable to the no-drive results for the effective hopping strength $J_{eff}=\epsilon/2$ at least for our accessible system size. With the growing belief of the absence of measurement-induced phase transition in the no-drive scenario Starchl et al. (2024b); Müller et al. (2025), it is tempting to conclude that in the high-frequency limit, maybe we don’t have any measurement-induced transition in the thermodynamic limit. No-drive results are presented in the Appendix D for our protocol. It is essential to note that the numerical data are not conclusive enough to confirm the absence of measurement-induced transitions in non-interacting lattice systems like ours. Moreover, the continuum models (which were primarily used as an analytical tool to predict the absence of transition) of Dirac fermions do not guarantee to describe the dynamics of monitored fermions on a tight-binding lattice, mostly because of the lack of energy conservation in such non-unitary dynamics Müller et al. (2025). On the other hand, we find that increasing $T$ tends to promote entanglement growth, and for reasonably large $T$ (in our manuscript, we primarily focus on $T=5$), the data suggest that the measurement-induced transition belonging to the BKT universality class may prevail even in the thermodynamic limit. However, in the absence of any concrete theoretical prediction, it is almost impossible to rule out the possibility that for extremely large system sizes, the steady-state entanglement data will keep showing area law scaling, and whatever we are observing in numerics is just a finite-size crossover between the logarithmic-area law phase. Even in this case, it is safe to say that with increasing $T$, the finite size crossover length scale will certainly increase, and for any reasonably large $T$, such system sizes will be almost inaccessible for any practical purposes Koh et al. (2023a). We also investigate a situation where we take a reasonably large $\gamma$ for which the $T=5$ data of the effective local central charge shows bending in the large subsystem size regime, and then we increase $T$. As expected, we could reduce the bending rate with increasing $T$, but were unable to achieve a scenario where the effective local central charge remains constant. Hence, the analysis does not allow us to conclude whether there is a finite critical driving period $T^{*}$ above which the bending completely stops. The absence of bending in the effective central charge data implies a critical phase. Hence, we also can’t rule out the possibility that beyond a specific critical measurement strength, the system will always remain in the area law phase for any finite driving period.
In this context, we also want to point out the many-body localized (MBL) systems. For a typical Heisenberg model, it was initially believed that the MBL transition was at disorder strength $W\simeq 3.5$ Luitz et al. (2015b). Then, there was a debate regarding the survival of the MBL phase in the thermodynamic limit Šuntajs et al. (2020a), and now it is believed that at least for $W<16$, there is no ”true” MBL transition in the thermodynamic limit Morningstar et al. (2022); Sels (2022), however, the most of the cold-atom experiments only can simulate the Heisenberg type model up to $L=20$ spins, and hence, for all practical purposes, the experiments only can probe that thermodynamically pre-thermal MBL phase, where dynamics is extremely slow Abanin et al. (2019). Hence, even if it ultimately turns out that the measurement-induced transition in driven non-interacting systems is not a true thermodynamic transition, but rather a finite-size crossover, at least it is clear that the crossover length scale increases with increasing driving period, hence, even such crossover will be viewed as effectively measurement induced transition in almost all experimental platforms Noel et al. (2022); Koh et al. (2023b); exp (2023).
@@
Moreover, we have investigated a drive protocol in which the hopping amplitudes are varied completely symmetrically around zero in the form of square-pulse and sinusoidal drives and found that such a drive can’t cause a phase transition independent of any frequency regime; it always promotes an area-law entanglement phase in the thermodynamic limit. This result is a bit surprising, especially given that disorder-induced similar-driven systems tend to favor de-localization in the low-frequency regime Sarkar et al. (2021). Our results suggest several interesting directions for future work. The role of disorders, interactions in the monitored Floquet system, and their relationship to Floquet-MBL Sierant et al. (2023) remain a challenging problem. It would be interesting if one could extend the analysis presented in Ref. Poboiko et al. (2023); Starchl et al. (2024b) to driven systems like ours in future studies.